# Vintix: Action Model via In-Context Reinforcement Learning

### v1

@@ Appendix F Dataset Size and Metadata
| Task | Trajectory Length | Number of trajectories | Mean Episode Length | State space shape | Action space shape | Reward Scaling |
| Hand-Block-Stack | 250000 | 10 | 248.6 | 1.0 | ||
| Hand-Bottle-Cap | 135000 | 10 | 123.7 | 1.0 | ||
| Hand-Catch-Abreast | 165000 | 10 | 98.5 | 1.0 | ||
| Hand-Catch-Over-2-Underarm | 100000 | 10 | 55.6 | 1.0 | ||
| Hand-Catch-Underarm | 100000 | 10 | 64.9 | 1.0 | ||
| Hand-Door-Close-Inward | 100000 | 10 | 23.8 | 1.0 | ||
| Hand-Grasp-And-Place | 500000 | 8 | 332.6 | 1.0 | ||
| Hand-Lift-Underarm | 500000 | 10 | 455.0 | 1.0 | ||
| Hand-Pen | 135000 | 10 | 120.5 | 1.0 | ||
| Hand-Push-Block | 135000 | 10 | 123.2 | 1.0 | ||
| Hand-Reorientation | 600000 | 7 | 463.9 | 1.0 | ||
| Hand-Scissors | 175000 | 10 | 149.0 | 1.0 | ||
| Hand-Swing-Cup | 320000 | 10 | 299.0 | 1.0 | ||
| Hand-Switch | 135000 | 10 | 124.0 | 1.0 | ||
| Hand-Two-Catch-Underarm | 100000 | 10 | 65.1 | 1.0 | ||
| Hand-Door-Close-Outward | 265000 | 10 | 249.0 | 1.0 | ||
| Hand-Door-Open-Inward | 265000 | 10 | 249.0 | 1.0 | ||
| Hand-Door-Open-Outward | 250000 | 10 | 249.0 | 1.0 | ||
| Hand-Kettle | 135000 | 10 | 124.0 | 1.0 | ||
@@
| Task | Trajectory Length | Number of trajectories | Mean Episode Length | State space shape | Action space shape | Reward Scaling |
| Ant-v4 | 1000000 | 10 | 524.1 | 0.1 | ||
| HalfCheetah-v4 | 1000000 | 5 | 1000 | 0.1 | ||
| Hopper-v4 | 1000000 | 10 | 227.3 | 0.1 | ||
| Humanoid-v4 | 1000000 | 15 | 207.7 | 0.1 | ||
| HumanoidStandup-v4 | 1000000 | 15 | 1000 | 0.01 | ||
| InvertedDoublePendulum-v4 | 1000000 | 10 | 55.7 | 0.1 | ||
| InvertedPendulum-v4 | 1000000 | 10 | 68.2 | 0.1 | ||
| Pusher-v5 | 1000000 | 5 | 100 | 0.1 | ||
| Reacher-v4 | 1000000 | 5 | 50 | 0.1 | ||
| Swimmer-v4 | 1000000 | 5 | 1000 | 1.0 | ||
@@
| Task | Trajectory Length | Number of trajectories | Mean Episode Length | State space shape | Action space shape | Reward Scaling |
| IB-p-0 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-5 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-10 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-15 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-20 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-25 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-30 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-35 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-40 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-45 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-50 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-55 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-60 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-65 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-70 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-75 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-80 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-85 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-90 | 100000 | 15 | 250 | 1.0 | ||
| IB-p-95 | 100000 | 15 | 250 | 1.0 | ||
@@
Appendix G Task-Level Performance
| Task | Random Score | Expert Score | Vintix (Normalized) |
| assembly | 45,1 3,3 | 297,9 24,8 | 1.04 0.10 |
| basketball | 3 1,3 | 558,4 69,3 | 1.02 0.11 |
| bin-picking | 2 0,6 | 424 105,7 | 0.01 0.01 |
| box-close | 78,7 11,6 | 520 122,6 | -0.04 0.03 |
| button-press | 31,3 4,3 | 642,9 13 | 0.97 0.07 |
| button-press-topdown | 31,6 9,8 | 483,7 41,7 | 0.94 0.14 |
| button-press-topdown-wall | 30,8 9 | 497,9 31,2 | 0.97 0.07 |
| button-press-wall | 10,3 4,4 | 675,9 15,2 | 1.00 0.04 |
| coffee-button | 33,1 7,8 | 731,4 29,1 | 1.00 0.06 |
| coffee-pull | 4,3 0,5 | 294 103,6 | 0.07 0.20 |
| coffee-push | 4,3 0,7 | 564,8 106,7 | 1.01 0.20 |
| dial-turn | 35,8 42,8 | 788,4 82,4 | 0.99 0.07 |
| disassemble | 40,6 6,8 | 527,3 64,6 | 1.00 0.12 |
| door-close | 5,5 1,4 | 541,9 25,6 | 1.02 0.05 |
| door-lock | 115,7 30,1 | 805,5 42,7 | 0.35 0.36 |
| door-open | 58,9 10,1 | 587,6 18,4 | 0.99 0.05 |
| door-unlock | 99,8 17,8 | 803,1 19,6 | 0.21 0.26 |
| drawer-close | 147,4 291,1 | 868,4 3,3 | 0.96 0.01 |
| drawer-open | 126,3 23,3 | 492,9 2,5 | 1.00 0.01 |
| faucet-close | 260,9 26,4 | 755,4 20,8 | 0.99 0.03 |
| faucet-open | 242,7 24,5 | 746,9 12,9 | 0.97 0.03 |
| hammer | 93,9 9,7 | 687,6 49,9 | 0.97 0.12 |
| hand-insert | 2,7 0,8 | 738,2 36,6 | 0.19 0.20 |
| handle-press | 73,6 76,1 | 812,8 170,8 | 1.00 0.15 |
| handle-press-side | 59,4 26,8 | 860,3 32,8 | 0.99 0.02 |
| handle-pull | 9,5 6,5 | 703,3 15,2 | 1.00 0.02 |
| handle-pull-side | 2,1 0,5 | 495,3 49,3 | 1.00 0.11 |
| lever-pull | 66,9 16,9 | 575,5 67,9 | 0.96 0.16 |
| peg-insert-side | 1,8 0,3 | 330,1 159,4 | 1.01 0.40 |
| peg-unplug-side | 4,6 2,5 | 528,8 87,1 | 0.75 0.34 |
| pick-out-of-hole | 8,6 49,9 | 401,1 105,8 | 0.91 0.20 |
| pick-place | 1,5 0,9 | 423,6 97,5 | 0.98 0.22 |
| pick-place-wall | 0 0 | 515,4 162,5 | 1.04 0.16 |
| plate-slide | 76,5 12 | 531,1 152 | 0.99 0.34 |
| plate-slide-back | 33,8 9,5 | 719,5 88 | 0.99 0.16 |
| plate-slide-back-side | 35,9 13 | 727,5 69,5 | 0.99 0.09 |
| plate-slide-side | 22,6 14 | 693,6 82,4 | 0.83 0.22 |
| push | 5,8 2,6 | 747,7 61,5 | 0.97 0.14 |
| push-back | 1,2 0,1 | 396,2 109,3 | 0.95 0.29 |
| push-wall | 6,5 4 | 749,7 11,7 | 0.99 0.02 |
| reach | 154 52,8 | 679,6 131,8 | 1.01 0.26 |
| reach-wall | 149 35,7 | 748,2 102,1 | 0.98 0.17 |
| shelf-place | 0 0 | 266,5 30 | 1.01 0.11 |
| soccer | 5,3 3,9 | 509,3 136,1 | 0.80 0.35 |
| stick-pull | 2,8 1,4 | 529,2 20,9 | 0.92 0.16 |
| stick-push | 2,9 1,1 | 626 46,9 | 0.90 0.28 |
| sweep | 12,4 8,5 | 518,4 48 | 0.94 0.12 |
| sweep-into | 16,9 14,3 | 799 14,9 | 1.00 0.02 |
| window-close | 58,2 15 | 591,1 38,6 | 0.95 0.17 |
@@ Appendix G Task-Level Performance
| Task | Random Score | Expert Score | Vintix (Normalized) |
| Ant-v4 | -459,2 824,1 | 6368,2 593,8 | 0,98 0,10 |
| HalfCheetah-v4 | -299,8 74,4 | 7782,8 109,2 | 0,93 0,03 |
| Hopper-v4 | 16,2 8,4 | 3237,8 707,8 | 0,86 0,19 |
| Humanoid-v4 | 116,5 31,6 | 7527,5 38,8 | 0,97 0,00 |
| HumanoidStandup-v4 | 37285,5 3178 | 300990,1 47970,1 | 1,02 0,21 |
| InvertedDoublePendulum-v4 | 56,2 16,1 | 9332,4 498,1 | 0,65 0,47 |
| InvertedPendulum-v4 | 5,6 2,1 | 1000 0 | 1,00 0,00 |
| Pusher-v5 | -151,9 8 | -40,1 7 | 1,02 0,08 |
| Reacher-v4 | -41,7 3,4 | -5,6 2,6 | 0,98 0,07 |
| Swimmer-v4 | 3 11,2 | 95,5 3,6 | 0,98 0,06 |
@@
| Task | Random Score | Expert Score | Vintix (Normalized) |
| shadowhandblockstack | 95,6 16 | 285 41,8 | 1.17 0.23 |
| shadowhandbottlecap | 110,6 17,6 | 399,9 57,6 | 0.81 0.25 |
| shadowhandcatchabreast | 1,1 0,6 | 65,6 14,2 | 0.17 0.32 |
| shadowhandcatchover2underarm | 4,9 0,7 | 34,2 6,1 | 0.92 0.24 |
| shadowhandcatchunderarm | 1,7 0,6 | 25,1 6,1 | 0.72 0.39 |
| shadowhanddoorcloseinward | 1,2 0,5 | 8,8 0,2 | 0.36 0.02 |
| shadowhanddoorcloseoutward | 941,2 43,8 | 1377,5 15,7 | -1.27 0.01 |
| shadowhanddooropeninward | -4,4 41,3 | 409,9 3,2 | 0.29 0.02 |
| shadowhanddooropenoutward | 20,4 40,9 | 617,1 4,4 | 0.13 0.02 |
| shadowhandgraspandplace | 6,8 1,6 | 498,1 51,1 | 0.97 0.18 |
| shadowhandkettle | -191,9 14,8 | 54,5 4,2 | -0.99 0.00 |
| shadowhandliftunderarm | -42,7 8,4 | 404 10,1 | 0.95 0.03 |
| shadowhandover | 2,7 0,7 | 34,6 5,8 | 0.95 0.03 |
| shadowhandpen | 4,5 2,4 | 186,3 19,5 | 0.52 0.44 |
| shadowhandpushblock | 224,7 66,5 | 457,4 3,6 | 0.98 0.01 |
| shadowhandreorientation | 127,1 434,9 | 3040,1 1986,8 | 0.89 0.66 |
| shadowhandscissors | -23,3 16,8 | 735,6 24,8 | 1.03 0.01 |
| shadowhandswingcup | -414,1 27,9 | 3937,9 601,1 | 0.95 0.13 |
| shadowhandswitch | 50,4 12,5 | 281,3 1,1 | 0.95 0.01 |
@@
| Task | Random Score | Expert Score | Vintix (Normalized) |
| industrial-benchmark-0-v1 | -348,9 32,7 | -180,7 3,1 | 0.94 0.13 |
| industrial-benchmark-5-v1 | -379,7 59 | -193,8 2,3 | 1.00 0.02 |
| industrial-benchmark-10-v1 | -395,4 68,9 | -215,1 2 | 1.01 0.01 |
| industrial-benchmark-15-v1 | -424,4 83 | -229,8 4 | 1.01 0.02 |
| industrial-benchmark-20-v1 | -455,8 88,2 | -249,4 2,2 | 0.95 0.11 |
| industrial-benchmark-25-v1 | -453,6 78,8 | -272,4 5,5 | 0.95 0.11 |
| industrial-benchmark-30-v1 | -480,3 76,6 | -288,3 5,4 | 0.90 0.10 |
| industrial-benchmark-35-v1 | -508,5 87 | -314,1 6,7 | 1.00 0.03 |
| industrial-benchmark-40-v1 | -515,8 77,4 | -337,8 8,8 | 0.99 0.05 |
| industrial-benchmark-45-v1 | -543,3 85,1 | -360,9 7,4 | 0.97 0.04 |
| industrial-benchmark-50-v1 | -574,9 69,8 | -377,9 7 | 0.91 0.09 |
| industrial-benchmark-55-v1 | -597,6 69,9 | -402,1 6 | 0.99 0.01 |
| industrial-benchmark-60-v1 | -625,2 83,4 | -430,3 4,8 | 0.98 0.02 |
| industrial-benchmark-65-v1 | -649,6 62,2 | -449,8 4,1 | 0.86 0.04 |
| industrial-benchmark-70-v1 | -691,7 87 | -471,1 4,2 | 0.95 0.03 |
| industrial-benchmark-75-v1 | -717,2 72,3 | -474,3 3,2 | 0.99 0.01 |
| industrial-benchmark-80-v1 | -757,8 105,9 | -485,4 3,1 | 0.96 0.01 |
| industrial-benchmark-85-v1 | -812,8 154,9 | -507,6 2,9 | 0.98 0.01 |
| industrial-benchmark-90-v1 | -846,4 132,7 | -522 3,6 | 0.97 0.01 |
| industrial-benchmark-95-v1 | -895,5 146,5 | -545,8 3,3 | 0.97 0.01 |

### v2

@@ Appendix G Dataset Size and Metadata
| Hand-Block-Stack | 250000 | 10 | 248.6 | 1.0 | |||||||||||||||
| Hand-Bottle-Cap | 135000 | 10 | 123.7 | 1.0 | |||||||||||||||
| Hand-Catch-Abreast | 165000 | 10 | 98.5 | 1.0 | |||||||||||||||
| Hand-Catch-Over-2-Underarm | 100000 | 10 | 55.6 | 1.0 | |||||||||||||||
| Hand-Catch-Underarm | 100000 | 10 | 64.9 | 1.0 | |||||||||||||||
| Hand-Door-Close-Inward | 100000 | 10 | 23.8 | 1.0 | |||||||||||||||
| Hand-Grasp-And-Place | 500000 | 8 | 332.6 | 1.0 | |||||||||||||||
| Hand-Lift-Underarm | 500000 | 10 | 455.0 | 1.0 | |||||||||||||||
| Hand-Pen | 135000 | 10 | 120.5 | 1.0 | |||||||||||||||
| Hand-Push-Block | 135000 | 10 | 123.2 | 1.0 | |||||||||||||||
| Hand-Reorientation | 600000 | 7 | 463.9 | 1.0 | |||||||||||||||
| Hand-Scissors | 175000 | 10 | 149.0 | 1.0 | |||||||||||||||
| Hand-Swing-Cup | 320000 | 10 | 299.0 | 1.0 | |||||||||||||||
| Hand-Switch | 135000 | 10 | 124.0 | 1.0 | |||||||||||||||
| Hand-Two-Catch-Underarm | 100000 | 10 | 65.1 | 1.0 | |||||||||||||||
| Hand-Door-Close-Outward | 265000 | 10 | 249.0 | 1.0 | |||||||||||||||
| Hand-Door-Open-Inward | 265000 | 10 | 249.0 | 1.0 | |||||||||||||||
| Hand-Door-Open-Outward | 250000 | 10 | 249.0 | 1.0 | |||||||||||||||
| Hand-Kettle | 135000 | 10 | 124.0 | 1.0 | |||||||||||||||
@@
| Ant-v4 | 1000000 | 10 | 524.1 | 0.1 | |||||||||||||||
| HalfCheetah-v4 | 1000000 | 5 | 1000 | 0.1 | |||||||||||||||
| Hopper-v4 | 1000000 | 10 | 227.3 | 0.1 | |||||||||||||||
| Humanoid-v4 | 1000000 | 15 | 207.7 | 0.1 | |||||||||||||||
| HumanoidStandup-v4 | 1000000 | 15 | 1000 | 0.01 | |||||||||||||||
| InvertedDoublePendulum-v4 | 1000000 | 10 | 55.7 | 0.1 | |||||||||||||||
| InvertedPendulum-v4 | 1000000 | 10 | 68.2 | 0.1 | |||||||||||||||
| Pusher-v5 | 1000000 | 5 | 100 | 0.1 | |||||||||||||||
| Reacher-v4 | 1000000 | 5 | 50 | 0.1 | |||||||||||||||
| Swimmer-v4 | 1000000 | 5 | 1000 | 1.0 | |||||||||||||||
@@
| IB-p-0 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-5 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-10 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-15 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-20 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-25 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-30 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-35 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-40 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-45 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-50 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-55 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-60 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-65 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-70 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-75 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-80 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-85 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-90 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
| IB-p-95 | 100000 | 15 | 250 | 1.0 | |||||||||||||||
@@
Appendix H Task-Level Performance
| assembly | 45,1 3,3 | 297,9 24,8 | 1.04 0.10 | |||||||
| basketball | 3 1,3 | 558,4 69,3 | 1.02 0.11 | |||||||
| bin-picking | 2 0,6 | 424 105,7 | 0.01 0.01 | |||||||
| box-close | 78,7 11,6 | 520 122,6 | -0.04 0.03 | |||||||
| button-press | 31,3 4,3 | 642,9 13 | 0.97 0.07 | |||||||
| button-press-topdown | 31,6 9,8 | 483,7 41,7 | 0.94 0.14 | |||||||
| button-press-topdown-wall | 30,8 9 | 497,9 31,2 | 0.97 0.07 | |||||||
| button-press-wall | 10,3 4,4 | 675,9 15,2 | 1.00 0.04 | |||||||
| coffee-button | 33,1 7,8 | 731,4 29,1 | 1.00 0.06 | |||||||
| coffee-pull | 4,3 0,5 | 294 103,6 | 0.07 0.20 | |||||||
| coffee-push | 4,3 0,7 | 564,8 106,7 | 1.01 0.20 | |||||||
| dial-turn | 35,8 42,8 | 788,4 82,4 | 0.99 0.07 | |||||||
| disassemble | 40,6 6,8 | 527,3 64,6 | 1.00 0.12 | |||||||
| door-close | 5,5 1,4 | 541,9 25,6 | 1.02 0.05 | |||||||
| door-lock | 115,7 30,1 | 805,5 42,7 | 0.35 0.36 | |||||||
| door-open | 58,9 10,1 | 587,6 18,4 | 0.99 0.05 | |||||||
| door-unlock | 99,8 17,8 | 803,1 19,6 | 0.21 0.26 | |||||||
| drawer-close | 147,4 291,1 | 868,4 3,3 | 0.96 0.01 | |||||||
| drawer-open | 126,3 23,3 | 492,9 2,5 | 1.00 0.01 | |||||||
| faucet-close | 260,9 26,4 | 755,4 20,8 | 0.99 0.03 | |||||||
| faucet-open | 242,7 24,5 | 746,9 12,9 | 0.97 0.03 | |||||||
| hammer | 93,9 9,7 | 687,6 49,9 | 0.97 0.12 | |||||||
| hand-insert | 2,7 0,8 | 738,2 36,6 | 0.19 0.20 | |||||||
| handle-press | 73,6 76,1 | 812,8 170,8 | 1.00 0.15 | |||||||
| handle-press-side | 59,4 26,8 | 860,3 32,8 | 0.99 0.02 | |||||||
| handle-pull | 9,5 6,5 | 703,3 15,2 | 1.00 0.02 | |||||||
| handle-pull-side | 2,1 0,5 | 495,3 49,3 | 1.00 0.11 | |||||||
| lever-pull | 66,9 16,9 | 575,5 67,9 | 0.96 0.16 | |||||||
| peg-insert-side | 1,8 0,3 | 330,1 159,4 | 1.01 0.40 | |||||||
| peg-unplug-side | 4,6 2,5 | 528,8 87,1 | 0.75 0.34 | |||||||
| pick-out-of-hole | 8,6 49,9 | 401,1 105,8 | 0.91 0.20 | |||||||
| pick-place | 1,5 0,9 | 423,6 97,5 | 0.98 0.22 | |||||||
| pick-place-wall | 0 0 | 515,4 162,5 | 1.04 0.16 | |||||||
| plate-slide | 76,5 12 | 531,1 152 | 0.99 0.34 | |||||||
| plate-slide-back | 33,8 9,5 | 719,5 88 | 0.99 0.16 | |||||||
| plate-slide-back-side | 35,9 13 | 727,5 69,5 | 0.99 0.09 | |||||||
| plate-slide-side | 22,6 14 | 693,6 82,4 | 0.83 0.22 | |||||||
| push | 5,8 2,6 | 747,7 61,5 | 0.97 0.14 | |||||||
| push-back | 1,2 0,1 | 396,2 109,3 | 0.95 0.29 | |||||||
| push-wall | 6,5 4 | 749,7 11,7 | 0.99 0.02 | |||||||
| reach | 154 52,8 | 679,6 131,8 | 1.01 0.26 | |||||||
| reach-wall | 149 35,7 | 748,2 102,1 | 0.98 0.17 | |||||||
| shelf-place | 0 0 | 266,5 30 | 1.01 0.11 | |||||||
| soccer | 5,3 3,9 | 509,3 136,1 | 0.80 0.35 | |||||||
| stick-pull | 2,8 1,4 | 529,2 20,9 | 0.92 0.16 | |||||||
| stick-push | 2,9 1,1 | 626 46,9 | 0.90 0.28 | |||||||
| sweep | 12,4 8,5 | 518,4 48 | 0.94 0.12 | |||||||
| sweep-into | 16,9 14,3 | 799 14,9 | 1.00 0.02 | |||||||
| window-close | 58,2 15 | 591,1 38,6 | 0.95 0.17 | |||||||
@@ Appendix H Task-Level Performance
| Ant-v4 | -459,2 824,1 | 6368,2 593,8 | 0,98 0,10 | |||||||
| HalfCheetah-v4 | -299,8 74,4 | 7782,8 109,2 | 0,93 0,03 | |||||||
| Hopper-v4 | 16,2 8,4 | 3237,8 707,8 | 0,86 0,19 | |||||||
| Humanoid-v4 | 116,5 31,6 | 7527,5 38,8 | 0,97 0,00 | |||||||
| HumanoidStandup-v4 | 37285,5 3178 | 300990,1 47970,1 | 1,02 0,21 | |||||||
| InvertedDoublePendulum-v4 | 56,2 16,1 | 9332,4 498,1 | 0,65 0,47 | |||||||
| InvertedPendulum-v4 | 5,6 2,1 | 1000 0 | 1,00 0,00 | |||||||
| Pusher-v5 | -151,9 8 | -40,1 7 | 1,02 0,08 | |||||||
| Reacher-v4 | -41,7 3,4 | -5,6 2,6 | 0,98 0,07 | |||||||
| Swimmer-v4 | 3 11,2 | 95,5 3,6 | 0,98 0,06 | |||||||
@@
| shadowhandblockstack | 95,6 16 | 285 41,8 | 1.17 0.23 | |||||||
| shadowhandbottlecap | 110,6 17,6 | 399,9 57,6 | 0.81 0.25 | |||||||
| shadowhandcatchabreast | 1,1 0,6 | 65,6 14,2 | 0.17 0.32 | |||||||
| shadowhandcatchover2underarm | 4,9 0,7 | 34,2 6,1 | 0.92 0.24 | |||||||
| shadowhandcatchunderarm | 1,7 0,6 | 25,1 6,1 | 0.72 0.39 | |||||||
| shadowhanddoorcloseinward | 1,2 0,5 | 8,8 0,2 | 0.36 0.02 | |||||||
| shadowhanddoorcloseoutward | 941,2 43,8 | 1377,5 15,7 | -1.27 0.01 | |||||||
| shadowhanddooropeninward | -4,4 41,3 | 409,9 3,2 | 0.29 0.02 | |||||||
| shadowhanddooropenoutward | 20,4 40,9 | 617,1 4,4 | 0.13 0.02 | |||||||
| shadowhandgraspandplace | 6,8 1,6 | 498,1 51,1 | 0.97 0.18 | |||||||
| shadowhandkettle | -191,9 14,8 | 54,5 4,2 | -0.99 0.00 | |||||||
| shadowhandliftunderarm | -42,7 8,4 | 404 10,1 | 0.95 0.03 | |||||||
| shadowhandover | 2,7 0,7 | 34,6 5,8 | 0.95 0.03 | |||||||
| shadowhandpen | 4,5 2,4 | 186,3 19,5 | 0.52 0.44 | |||||||
| shadowhandpushblock | 224,7 66,5 | 457,4 3,6 | 0.98 0.01 | |||||||
| shadowhandreorientation | 127,1 434,9 | 3040,1 1986,8 | 0.89 0.66 | |||||||
| shadowhandscissors | -23,3 16,8 | 735,6 24,8 | 1.03 0.01 | |||||||
| shadowhandswingcup | -414,1 27,9 | 3937,9 601,1 | 0.95 0.13 | |||||||
| shadowhandswitch | 50,4 12,5 | 281,3 1,1 | 0.95 0.01 | |||||||
@@
| industrial-benchmark-0-v1 | -348,9 32,7 | -180,7 3,1 | 0.94 0.13 | |||||||
| industrial-benchmark-5-v1 | -379,7 59 | -193,8 2,3 | 1.00 0.02 | |||||||
| industrial-benchmark-10-v1 | -395,4 68,9 | -215,1 2 | 1.01 0.01 | |||||||
| industrial-benchmark-15-v1 | -424,4 83 | -229,8 4 | 1.01 0.02 | |||||||
| industrial-benchmark-20-v1 | -455,8 88,2 | -249,4 2,2 | 0.95 0.11 | |||||||
| industrial-benchmark-25-v1 | -453,6 78,8 | -272,4 5,5 | 0.95 0.11 | |||||||
| industrial-benchmark-30-v1 | -480,3 76,6 | -288,3 5,4 | 0.90 0.10 | |||||||
| industrial-benchmark-35-v1 | -508,5 87 | -314,1 6,7 | 1.00 0.03 | |||||||
| industrial-benchmark-40-v1 | -515,8 77,4 | -337,8 8,8 | 0.99 0.05 | |||||||
| industrial-benchmark-45-v1 | -543,3 85,1 | -360,9 7,4 | 0.97 0.04 | |||||||
| industrial-benchmark-50-v1 | -574,9 69,8 | -377,9 7 | 0.91 0.09 | |||||||
| industrial-benchmark-55-v1 | -597,6 69,9 | -402,1 6 | 0.99 0.01 | |||||||
| industrial-benchmark-60-v1 | -625,2 83,4 | -430,3 4,8 | 0.98 0.02 | |||||||
| industrial-benchmark-65-v1 | -649,6 62,2 | -449,8 4,1 | 0.86 0.04 | |||||||
| industrial-benchmark-70-v1 | -691,7 87 | -471,1 4,2 | 0.95 0.03 | |||||||
| industrial-benchmark-75-v1 | -717,2 72,3 | -474,3 3,2 | 0.99 0.01 | |||||||
| industrial-benchmark-80-v1 | -757,8 105,9 | -485,4 3,1 | 0.96 0.01 | |||||||
| industrial-benchmark-85-v1 | -812,8 154,9 | -507,6 2,9 | 0.98 0.01 | |||||||
| industrial-benchmark-90-v1 | -846,4 132,7 | -522 3,6 | 0.97 0.01 | |||||||
| industrial-benchmark-95-v1 | -895,5 146,5 | -545,8 3,3 | 0.97 0.01 | |||||||
@@
Appendix I Comparison with other cross-domain agents
| Ant | 0.98 0.10 | 0.17 | 0.88 0.29 | ||||
| HalfCheetah | 0.93 0.03 | 0.34 | 0.89 0.03 | ||||
| Hopper | 0.86 0.19 | 0.12 | 0.76 0.21 | ||||
| Humanoid | 0.97 0.00 | 0.02 | 0.10 0.02 | ||||
| HumanoidStandup | 1.02 0.21 | 0.26 | 0.35 0.09 | ||||
| InvertedDoublePendulum | 0.65 0.47 | 0.02 | 0.93 0.14 | ||||
| InvertedPendulum | 1.00 0.00 | 0.06 | 0.24 0.05 | ||||
| Pusher | 1.02 0.08 | 0.90 | 1.00 0.05 | ||||
| Reacher | 0.98 0.07 | 0.90 | 0.99 0.06 | ||||
| Swimmer | 0.98 0.06 | 0.82 | 1.02 0.04 | ||||
| Walker2d | 1.00 0.02 | 0.05 | 0.95 0.18 |
| assembly | 1.04 0.10 | 0.83 | 0.96 0.17 | ||||
| basketball | 1.02 0.11 | 0.68 | -0.00 0.00 | ||||
| bin-picking | 0.01 0.01 | 0.67 | 0.47 0.52 | ||||
| box-close | -0.04 0.03 | 0.93 | 0.89 0.39 | ||||
| button-press | 0.97 0.07 | 0.62 | 0.86 0.30 | ||||
| button-press-topdown | 0.94 0.14 | 0.62 | 0.51 0.17 | ||||
| button-press-topdown-wall | 0.97 0.07 | 0.62 | 0.53 0.19 | ||||
| button-press-wall | 1.00 0.04 | 0.94 | 0.94 0.18 | ||||
| coffee-button | 1.00 0.06 | 0.84 | 0.38 0.41 | ||||
| coffee-pull | 0.07 0.20 | 0.62 | 0.14 0.27 | ||||
| coffee-push | 1.01 0.20 | 0.18 | 0.30 0.44 | ||||
| dial-turn | 0.99 0.07 | 0.83 | 0.95 0.16 | ||||
| disassemble | 1.00 0.12 | 2.24 | 0.17 3.91 | ||||
| door-close | 1.02 0.05 | 1.00 | 0.99 0.06 | ||||
| door-lock | 0.35 0.36 | 0.85 | 0.84 0.28 | ||||
| door-open | 0.99 0.05 | 0.98 | 0.99 0.10 | ||||
| door-unlock | 0.21 0.26 | 0.90 | 0.95 0.13 | ||||
| drawer-close | 0.96 0.01 | 1.00 | 0.64 0.30 | ||||
| drawer-open | 1.00 0.01 | 0.96 | 0.98 0.10 | ||||
| faucet-close | 0.99 0.03 | 0.53 | 0.23 0.18 | ||||
| faucet-open | 0.97 0.03 | 0.99 | 0.70 0.37 | ||||
| hammer | 0.97 0.12 | 0.95 | 0.96 0.15 | ||||
| hand-insert | 0.19 0.20 | 0.82 | 0.93 0.25 | ||||
| handle-press | 1.00 0.15 | 0.99 | 0.84 0.32 | ||||
| handle-press-side | 0.99 0.02 | 0.99 | 0.01 0.09 | ||||
| handle-pull | 1.00 0.02 | 0.48 | 0.83 0.25 | ||||
| handle-pull-side | 1.00 0.11 | 0.48 | 0.50 0.49 | ||||
| lever-pull | 0.96 0.16 | 0.19 | 0.40 0.43 | ||||
| peg-insert-side | 1.01 0.40 | 0.70 | 0.81 0.51 | ||||
| peg-unplug-side | 0.75 0.34 | 0.31 | 0.17 0.32 | ||||
| pick-out-of-hole | 0.91 0.20 | 0.01 | 0.00 0.00 | ||||
| pick-place | 0.98 0.22 | 0.99 | 0.32 0.48 | ||||
| pick-place-wall | 1.04 0.16 | 0.99 | 0.10 0.29 | ||||
| plate-slide | 0.99 0.34 | 1.00 | 0.90 0.42 | ||||
| plate-slide-back | 0.99 0.16 | 1.00 | 0.24 0.00 | ||||
| plate-slide-back-side | 0.99 0.09 | 1.00 | 0.96 0.17 | ||||
| plate-slide-side | 0.83 0.22 | 0.99 | 0.16 0.04 | ||||
| push | 0.97 0.14 | 0.84 | 0.94 0.21 | ||||
| push-back | 0.95 0.29 | 0.84 | 0.97 1.29 | ||||
| push-wall | 0.99 0.02 | 0.81 | 0.21 0.30 | ||||
| reach | 1.01 0.26 | 0.99 | 0.34 0.32 | ||||
| reach-wall | 0.98 0.17 | 0.99 | 0.81 0.37 | ||||
| shelf-place | 1.01 0.11 | 0.96 | 0.38 0.46 | ||||
| soccer | 0.80 0.35 | 0.61 | 0.77 0.44 | ||||
| stick-pull | 0.92 0.16 | 0.88 | 0.92 0.23 | ||||
| stick-push | 0.90 0.28 | 0.75 | 0.48 0.48 | ||||
| sweep | 0.94 0.12 | 0.91 | 0.01 0.04 | ||||
| sweep-into | 1.00 0.02 | 0.91 | 0.99 0.06 | ||||
| window-close | 0.95 0.17 | 0.90 | 0.67 0.39 | ||||
| window-open | 0.96 0.17 | 0.97 | 0.98 0.12 |